### v1

# Multimodal Crash Likelihood Prediction: A Complexity-Infused Approach Integrating Semantic, Contextual, and Driving Features

Abstract
Predicting crash likelihood in complex driving environments is essential for improving traffic safety and advancing autonomous driving. Previous studies have used statistical models and deep learning to predict crashes based on semantic, contextual, or driving features, but none have examined the combined influence of these factors, termed “roadway complexity” in this study. This paper introduces a two-stage framework that integrates roadway complexity features for crash prediction. In the first stage, an encoder extracts hidden contextual information from these features, generating complexity-infused features. The second stage uses both original and complexity-infused features to predict crash likelihood, achieving an accuracy of 87.98% with original features alone and 90.15% with the added complexity-infused features. Ablation studies confirm that a combination of semantic, driving, and contextual features yields the best results, which emphasize their role in capturing roadway complexity. Additionally, complexity index annotations generated by Large Language Models outperform those by Amazon Mechanical Turk, highlighting the potential of automated tools for accurate, scalable crash prediction systems.
1 Introduction
Figure 1: The model structure. The model takes raw images as input and generates semantic, driving, and contextual features from various sources to predict the crash likelihood as the output. It consists of an encoder that learns hidden features from the semantic, driving, and contextual data, which are infused with the complexity index. The prediction model then utilizes all the available features, including the complexity-infused features, to predict the crash likelihood. Example data is shown above each feature source.
Understanding factors contributing to crashes is essential for improving traffic safety and advancing autonomous vehicle design. A 2015 NHTSA report [20] found that 94% of the 5,470 crashes from 2005 to 2007 were due to human errors like inattention, distractions, excessive speed, and misjudgment.
Research has sought to predict crash frequency and rates based on environmental factors and driver behavior. Statistical models have linked crash occurrences to variables such as speed, traffic volume, weather, and road design [22, 19, 24, 6, 23], while recent studies have used deep learning to analyze real-time images and sensor data from vehicles and infrastructure to detect hazards [11, 8]. Though promising, these models rely on either imagery or behavioral data, without integrating both to predict crashes.
In this paper, we define roadway scene complexity as a combination of semantic information (e.g., number of objects) and contextual variation (e.g., road curvature, roadway type, weather). Scene complexity affects driver behavior, with factors like object density and road type impacting drivers’ situational awareness [18, 1]. Visual and environmental complexity, such as high traffic volumes or urban versus rural settings, also influences cognitive demands on drivers [10, 7, 15, 17].
Driving behavior, such as speed adjustments in response to poor visibility or narrow lanes, is also influenced by scene complexity. Speed and acceleration patterns adjust based on obstacles and conditions [12]. Integrating behavior data with scene information deepens our understanding of driver interactions with their environment, improving crash risk modeling. Thus, we define roadway complexity as the combination of scene complexity and driving behavior.
Extracting hidden context from this combined data is essential. Previous studies have shown that fusing situational and memory-based features [27], as well as road graph and motion history data [16], enhances situation awareness and motion prediction, respectively. Building on this, we incorporate feature fusion to capture both explicit and implicit features of roadway scenes.
To address the challenges of (1) studying roadway complexity more holistically by incorporating both imagery and behavioral data, and (2) investigating the direct and indirect relationships between roadway complexity and crashes, we propose a two-stage fusion prediction model framework. This framework learns hidden features from the fusion of scene-level semantic, driving-level, and contextual features and uses the combination of these features to estimate crash likelihood, as illustrated in Fig. 1. Our model, which is trained on a naturalistic dataset and historical crash data, consists of an encoder that captures the hidden context of the roadway complexity and a prediction module that uses these features to estimate crash likelihood. The contributions of this paper are:
  1. 1.
    We take a holistic approach to studying roadway complexity by considering both imagery information, i.e., specifically, scene-level semantic and contextual features, and driving behavioral data, and we empirically demonstrate the necessity of incorporating all these elements.
  2. 2.
    We argue that crashes can be directly and/or indirectly associated with roadway complexity. We introduce a two-stage framework that extracts hidden contextual features of roadway scene complexity by fusing scene-level, driving-level, and contextual data, and we investigate the predictive power of both direct and hidden features on crash likelihood.
  3. 3.
    We compare the complexity index annotations from Amazon Mechanical Turk and Large Language Models (LLMs) in terms of their capability to predict crash likelihood and find that LLMs-generated annotations consistently exhibited better predictive performance. This can enhance the development of real-time crash prediction systems and inspire the integration of automated annotation tools for improved accuracy and scalability.
2 Related Work
There has been extensive research on understanding roadway scenes using both Convolutional Neural Network (CNN)-based and transformer-based methods. Panoptic segmentation, which unifies semantic and instance segmentation, offers a holistic understanding by labeling all pixels and differentiating object instances, making it especially useful in structured environments like roadways [13]. The task has been successfully applied to driving scene datasets, such as CityScapes [2], which provides high-resolution images for semantic urban scene understanding. Recent advancements in transformer-based models have further improved performance on scene understanding tasks. Among these, OneFormer [9] has achieved state-of-the-art panoptic quality scores [13] while maintaining computational efficiency, making it suitable for real-time scene understanding in autonomous driving. Given its strong performance, the OneFormer model was utilized in this study to extract scene-level visual features from the driving environments.
Understanding the driving context is essential for both human drivers and autonomous systems, as it involves not only the perception of roadway scene elements but also can help interpret the interactions between these elements. Recent advancements in LLMs have demonstrated their potential to go beyond pixel-level interpretation, incorporating higher-level reasoning into driving tasks. LLMs have been applied to extract contextual information from driving environments by mapping raw sensory data (e.g., images, LiDAR) to higher-level contextual descriptions [21]. This is particularly useful for driving research, where understanding the “intent” of other road users, predicting possible hazards, or inferring rules of the road from visual cues can significantly enhance safety and performance [3, 26]. The application of LLMs indicates a leap forward as they can go beyond object detection to understand the context of the driving scenarios.
3 Methods
In this section, we outline the dataset that was used in the study and the methodology for generating the features used in the crash likelihood prediction model. We detail the architecture of the complexity-infused encoder and describe the structure of the final prediction model.
3.1 Dataset Overview
This study utilized the MIT-AVT dataset as input [5]. The dataset includes a variety of multimodal data sources, such as 20-second raw video clips from the forward-facing camera, 30 Hz CAN bus data, 30 Hz GPS data, and contextual metadata features, as described by Ding et al. [4].
For the purposes of this research, 500 video clips were selected from the original dataset, categorized as follows: 100 highway scenarios, 100 rural scenarios, 100 urban scenarios, 75 bridge scenarios, 75 overpass scenarios, and 75 crash hotspot scenarios. The methodology for generating crash hotspot data is detailed in Section 3.3.
From each selected video clip, frames were extracted based on a fixed-distance sampling method, with one frame captured for every 20 meters traveled. The extraction began with the first frame of each clip and continued until the distance to the subsequent frame was less than 20 meters. Overall, there are around 10,300 frames extracted.
3.2 Feature Generation
For each image frame, three sets of features were generated: semantic features derived from a computer vision algorithm, driving features extracted from the CAN bus data, and contextual features generated from the LLMs.
3.2.1 Semantic Features
The semantic features were generated using the OneFormer algorithm via Hugging Face’s API [^1] https://huggingface.co/docs/transformers/en/model_doc/oneformer. The model outputs pixel-level semantic classifications for various objects such as cars, pedestrians, bicycles, roads, traffic signs, sidewalks, buildings, vegetation, and sky. To better understand the effect of complexity on driver behavior, a lead-car region was defined [25]. An example of the OneFormer algorithm’s output is provided in Fig. 2.
Figure 2: The raw roadway scene image and OneFormer algorithm output. The lead-car region is highlighted in a green box.
For each image, the percentage of pixels corresponding to each class relative to the total frame size was calculated. Similarly, the percentage of pixels for each class within the lead-car region was computed. Additionally, the number of cars, pedestrians, buses, bicycles, and motorcycles was counted in both the full frame and the lead-car regions. This process resulted in 50 initial semantic features. After closer examination, features with minimal variability (more than 90% being 0) were removed, reducing the final set to 17 semantic features. The final set of features is coded as:
  1. 1.
    Full frame (10): ‘car_count’, ‘road’, ‘vegetation’, ‘sky’, ‘terrain’, ‘car’, ‘sidewalk’, ‘building’, ‘traffic_light’, ‘person’.
  2. 2.
    Lead car region (7): ‘lead_car_traffic_sign’, ‘lead_car_road’, ‘lead_car_vegetation’, ‘lead_car_sky’, ‘lead_car_car’, ‘lead_car_fence’, ‘lead_car_car_count’.
3.2.2 Driving Features
The driving features were extracted from the CAN bus data. For each frame and the corresponding 20-meter segment, the following 9 features were computed: current speed, mean speed, standard deviation of speed, mean longitudinal acceleration, standard deviation of longitudinal acceleration, minimum longitudinal acceleration, maximum longitudinal acceleration, raw deviation from the speed limit, and normalized deviation from the speed limit. The raw and normalized deviation from the speed limit are defined as follows:
$SpeedDev=Speed_{mean}-SpeedLimit$ (1)
${SpeedDev}_{normalized}=\frac{Speed_{mean}-SpeedLimit}{SpeedLimit}$ (2)
To obtain the speed limit information for each road segment, Microsoft Azure’s Get Search Address Reverse API  [^2] https://learn.microsoft.com/en-us/rest/api/maps/search/get-search-address-reverse?view=rest-maps-1.0&tabs=HTTP was utilized.
3.2.3 Contextual Features
The contextual features include road characteristics, which were generated using the GPT-4o-2024-08-06 model  [^3] https://platform.openai.com/docs/models/gpt-4o. The prompt used in this step is illustrated in Fig. 3. In the prompt, several questions were asked to gather contextual road characteristics, including information on weather conditions, road conditions, traffic conditions, visibility levels, time of day, road layout, road type, and lane width. Initially, each question was presented in an open-ended format to generate a pool of possible answers. The prompt was then refined based on these responses, providing predefined options to ensure consistency across the answers. The results were output in JSON format. The prompt was run three times on each image to ensure data quality. The answer agreed by the majority will be chosen as the final answer to the questions.
During analysis, these contextual features were one-hot encoded to convert them into numerical values for integration into the model. In total, this process generated 19 contextual features. As can be seen in Fig. 3, there were 5 features for weather, 2 for road condition, 2 for traffic condition, 1 for visibility, 2 for time of day, 2 for road layout, 3 for road type, and 2 for lane width.
Figure 3: The prompt used in collecting contextual features with GPT-4o model.
3.3 Ground Truth Generation
There are two ground truths in this study. As illustrated in Figure 1, for the encoder, the complexity index is collected to quantify the overall complexity level perceived by the driver in a given roadway scene. For the prediction, crash likelihood is computed and serves as the ground truth for estimating crash probability.
3.3.1 Complexity Index
The complexity index was generated from two sources: AI and humans. For AI, the GPT-4o-2024-08-06 model was used along with the contextual feature generation process, as shown in Fig. 3. In this approach, the model generated a complexity score on a scale from 0 to 10 to describe the complexity and demanding level of the roadway scenes.
The human-generated complexity indices relied on Amazon Mechanical Turk (MTurk) for annotations. The task was designed to assess the complexity level of roadway scenes. Workers were shown image frames and asked to rate the complexity of each scene on a scale from 1 to 10. Only workers with a high approval rating, at least 500 completed tasks, and residing in the US were selected. A pilot study was conducted with 500 images, where 10 workers annotated the same image. The results showed a relatively high level of agreement among workers. Based on this, in the official round, each scene was annotated by 3 workers, and the final complexity score was determined by averaging their responses.
3.3.2 Crash Likelihood
To generate the crash likelihood, historical crash data was obtained from the Massachusetts Department of Transportation’s IMPACT app  [^4] https://apps.impact.dot.state.ma.us/cdp/home. GPS data from crashes over a 5-year period (2018 to 2022) was aggregated. A Kernel Density Estimation (KDE) method was then applied to this GPS data to create a continuous scale representing crash likelihood. The Kernel Density Estimation is formulated as follows:
$Density=\frac{1}{(radius)^{2}}\sum[\frac{3}{\pi}\cdot(1-(\frac{{dist}_{i}}{radius})^{2})^{2}],$ (3)
for ${dist}_{i}<radius$, where $i=1,\dots,n$ are the input points. Only the points within the $radius$ distance of the $(x,y)$ location will be included. ${dist}_{i}$ is the distance between point $i$ and the $(x,y)$ location. In this study, a radius of 1000 meters was used.
After applying the KDE method, each location was assigned a corresponding crash likelihood. The GPS data from the MIT-AVT video clips was then mapped onto this heatmap so that each 20-meter segment is associated with an estimated crash likelihood. The resulting crash likelihood heatmap and the 500 video clips in the MIT-AVT dataset are shown in Fig. 4.
Figure 4: Crash likelihood heatmap (2018-2022) in Massachusetts, displayed in red, where darker colors indicate a higher crash likelihood. Five hundred video clips from the MIT-AVT dataset are marked in blue.
To keep the scale consistent with the Complexity Index, the crash likelihood is normalized and represented on a scale from 0 to 10. The 10,300 frames were categorized into three levels based on the calculated crash likelihood. As shown in Fig. 5, the crash likelihood distribution is skewed and unbalanced. There is a noticeable trend where crash likelihoods between 0 and 0.5 are highly frequent, decrease somewhat in the range of 0.5 to 2, and become increasingly rare beyond 2. Therefore, likelihoods between 0 and 0.5 were categorized as Low, those between 0.5 and 2 as Medium, and those between 2 and 10 as High. This classification resulted in approximately 3,600 images in the low category, 2,800 in the medium category, and 3,700 in the high category.
Figure 5: The distribution of crash likelihood.
3.4 Modeling
3.4.1 Complexity-infused Encoder
The complexity-infused encoder uses a fully connected neural network structure with either 16 or 32 hidden neurons. The input to the network is threefold: (1) the 17 semantic features, (2) the combination of the 17 semantic features and 9 driving features, or (3) the combination of all features—17 semantic features, 9 driving features, and 19 contextual features. The input variables are normalized to a 0-1 range to ensure consistency across features and improve the stability of the model during training. The output of the network is the complexity index, which is treated as either a continuous or categorical variable for data obtained from the LLM, and as a continuous variable for data obtained from MTurk. The Root Mean Square Error (RMSE) is used as the evaluation metric for the complexity index when treated as a continuous variable, while accuracy is used as the metric when the complexity index is treated as a categorical variable.
After the hidden layer, a ReLU activation function is applied. The features generated after the ReLU activation are referred to as the complexity-infused features.
3.4.2 Crash Prediction Model
After generating the complexity-infused features from the encoder, they were used to predict the level of crash likelihood in combination with the corresponding input feature sets. For example, if the complexity-infused features were trained on only semantic features, the input for the crash prediction model would consist of both the complexity-infused and semantic features. Similarly, if the complexity-infused features were trained on all available features, the input to the crash prediction model would include the complexity-infused, semantic, driving, and contextual features.
To train the crash prediction model, various algorithms were tested, including Random Forest (RF), Gradient-Boosted Decision Trees (GBDT), K-Nearest Neighbors (KNN), and fully connected neural networks (NN). The NN model used in this step is a fully connected neural network consisting of seven linear layers. Similar to the encoder described in Section 3.4.1, the input variables are normalized to a 0-1 range. The dataset was split into 70% for training and 30% for testing, and model performance was evaluated using accuracy as the primary metric. To ensure consistency, the dataset was split in the same way as it was for the encoder.
4 Experiments and Results
4.1 Complexity-infused Encoder
Implementation. The complexity-infused encoder was implemented with different settings, depending on the input feature sets and the type of output variable, i.e., the complexity index. When the complexity index was treated as a continuous variable, the Mean Squared Error (MSE) was used as the loss function. When the index was treated as a categorical variable, Cross Entropy (CE) was chosen as the loss function. Further details of the model training configurations can be found in the Appendix.
Results. The model performance of the encoder is presented in Table 1, showing the model performance on the complexity index obtained from the LLMs. As shown in the table, the three model settings exhibited similar performance across different input feature sets. However, several insights can be drawn from the results: first, the model with 32 hidden neurons and a continuous output produced the lowest RMSE overall. Second, models with more comprehensive input feature sets, which incorporated semantic, driving, and contextual features, consistently outperformed those with fewer feature sets. This pattern was observed across all three models.
Model Settings Input Features
RMSE/CE
on Train $\downarrow$
RMSE/CE
on Test $\downarrow$
32 neurons, 1-d cont. output Semantic $1.10$ $1.10$
Semantic + Driving $1.05$ $1.06$
Sem. + Driv. + Cont. $0.84$ $0.86$
16 neurons, 1-d cont. output Semantic $1.11$ $1.12$
Semantic + Driving $1.08$ $1.08$
Sem. + Driv. + Cont. $0.84$ $0.85$
32 neurons, 10-d cat. output Semantic $1.33$ $1.35$
Semantic + Driving $1.31$ $1.34$
Sem. + Driv. + Cont. 1.15.15.15.15 $1.18$
Table 1: Performance comparison of the complexity-infused encoder models with different input feature sets and hidden neurons. The results are reported in terms of RMSE for continuous outputs and Cross-Entropy for categorical outputs.
Since the encoder with 32 hidden neurons and a continuous output variable yielded the best results, the 32 hidden features from this model will be used as the complexity-infused features in Section 4.2.
4.2 Main Results of Crash Prediction Model
Implementation. When training the neural network for the crash prediction model, Cross Entropy was chosen as the loss function during training. The accuracy of the best-performing model will be reported in the subsequent section. Further details on the model training configurations can be found in the Appendix.
Results. The performance of the crash prediction model is shown in Table 2. The baseline model was trained using only the original feature sets, and the complexity-infused features were subsequently added to investigate any improvement in model performance. Additionally, for comparison, the performance of the model trained using only the complexity-infused features was evaluated, as well as the effect of adding the complexity index to the baseline model.
Input Features Model Baseline $\uparrow$ + Comp.-infused $\uparrow$ Difference Comp.-infused alone $\uparrow$ + Comp. Index $\uparrow$
Semantic RF 73.23 78.35 5.12 65.31 74.07
GBDT 68.42 72.49 4.07 60.97 68.30
KNN 66.09 70.86 4.77 63.85 67.49
NN 67.81 73.11 5.30 61.07 68.83
Semantic + Driving RF 83.56 86.32 2.76 67.85 84.07
GBDT 74.62 77.41 2.79 62.24 74.88
KNN 74.15 75.98 1.83 67.42 73.53
NN 75.60 77.95 2.35 66.08 72.45
Semantic + Driving + Contextual RF 87.98 90.15 2.17 78.49 87.35
GBDT 80.13 82.24 2.11 68.03 80.04
KNN 81.41 82.48 1.07 76.74 81.25
NN 80.08 88.78 8.70 73.01 84.53
Table 2: Performance comparison of different combinations of feature sets for crash likelihood prediction. The results are reported in terms of accuracy (%) on the test set. Upward arrows next to the values indicate that higher accuracy is preferable.
The results indicated that the Random Forest model consistently achieved the best performance across all combinations of input feature sets. Additionally, there was a clear trend of improved model performance as the number of input features increased, with the highest accuracy reaching 87.98%. Adding the complexity-infused features led to further improvements in prediction performance, with the highest accuracy being 90.15%, when using semantic + driving + contextual features. This was followed by the neural network model, which achieved an accuracy of $88.78\%$. To assess the significance of these improvements, McNemar’s test [14] was conducted for each model, and the results indicated that the improvements were statistically significant. The results suggested that incorporating complexity-infused features, particularly in models using a combination of semantic, driving, and contextual features, significantly enhances the accuracy of crash likelihood predictions.
For the best-performing model, which was trained on all available features, including complexity-infused features, using the Random Forest algorithm, the feature importance is shown in Fig. 6. The results indicate that, among the semantic features, the most important factors in classifying crash likelihood include the vegetation percentage, sky percentage, road percentage, car count, and car percentage, both in the full frame and lead-car region.
For driving features, key factors include the original speed, mean and standard deviation of speed, raw and normalized speed deviation from the speed limit, and the standard deviation of longitudinal acceleration. Additionally, four hidden complexity-infused features (labeled 0, 1, 21, and 30) were identified as important. Notably, none of the contextual features were ranked among the top 20 most important features in the crash prediction model.
Figure 6: The feature importance of the best-performing model.
Using the complexity-infused features alone or adding the 1-dimensional complexity index to the baseline models did not outperform the baseline models, except for the neural network model on Semantic + Driving + Contextual features. This suggests that simply adding the complexity index does not substantially improve the model’s prediction capability. However, the complexity-infused features were found to effectively enhance model performance. This indicates that the complexity index is closely associated with scene-level semantic features, driving behavior, and contextual characteristics, and that hidden information can be effectively extracted using neural network structures.
4.3 Ablation Studies
Several experiments are conducted to justify the usability, stability, and robustness of the proposed Complexity-infused prediction model.
4.3.1 Comparison of LLMs and MTurk Annotations
The encoder trained on the MTurk annotations utilized a 32-hidden neuron structure and treated the output as a continuous variable, since the annotations were generated by averaging the workers’ responses. The performance of the encoders trained on LLMs and MTurk data is shown in Table 3. It can be seen that the encoder using MTurk data performed slightly better than the one using LLMs data. However, for the encoders trained on the combination of semantic, driving, and contextual features, the performance of both encoders was very similar.
Annotation
Sources
Input Features
RMSE/CE
on Train $\downarrow$
RMSE/CE
on Test $\downarrow$
LLMs Semantic 1.10 1.10
Semantic+Driving 1.05 1.06
Sem. + Driv. + Cont. 0.84 0.86
MTurk Semantic 0.93 0.94
Semantic+Driving 0.91 0.93
Sem. + Driv. + Cont. 0.83 0.88
Table 3: Performance comparison of the complexity-infused encoder models using LLMs data or MTurk data. The results are reported in terms of RMSE. Downward arrows next to the values indicate that lower RMSE is preferable.
Next, the crash prediction model was trained using the complexity-infused features derived from the MTurk data. The results of the crash likelihood prediction models trained on both LLMs and MTurk data are shown in Table 4. The model using complexity-infused features derived from LLMs data consistently yielded higher accuracy than the one using MTurk data. This suggests that the LLMs-generated complexity indexes may capture relevant information more effectively for crash likelihood prediction.
Annotation
Sources
Input Features
Comp.-infused
+ Base $\uparrow$
Comp.-infused
alone $\uparrow$
LLMs Semantic 78.35 65.31
Semantic+Driving 86.32 67.85
Sem. + Driv. + Cont. 90.15 78.49
MTurk Semantic 71.01 67.72
Semantic+Driving 80.89 73.61
Sem. + Driv. + Cont. 86.60 80.06
Table 4: Performance comparison of the crash prediction models using complexity-infused features from LLMs data or MTurk data. The results are reported in terms of accuracy (%). Upward arrows next to the values indicate that higher accuracy is preferable.
It is also worth noting that the model using only the complexity-infused features from MTurk annotations outperformed the one based on LLMs annotations. When considering the encoder performance (Table 3), the model trained solely on complexity-infused features from MTurk data showed better prediction capability than the ones from LLMs data. However, when semantic, driving, and contextual features were added, the best-performing model was achieved using the LLMs data. This implies that while the complexity-infused features derived from MTurk annotations capture important information effectively in isolation, the LLMs-generated data provides a more robust integration of semantic, driving, and contextual features. As a result, the combination of LLMs data with these additional features enhances the model’s overall performance.
4.3.2 Comparisons of Other Encoders
Similar to the experiments conducted in Section 4.2, the crash prediction model was trained using complexity-infused features from the second and third encoders, as presented in Table 1. The results are shown in Table 5.
Encoder Input Features
Base +
Comp.-infused $\uparrow$
32 neurons, 1-d cont. output Semantic 78.35
Semantic + Driving 86.32
Sem. + Driv. + Cont. 90.15
16 neurons, 1-d cont. output Semantic 77.42
Semantic + Driving 86.24
Sem. + Driv. + Cont. 89.56
32 neurons, 10-d cat. output Semantic 75.07
Semantic + Driving 83.59
Sem. + Driv. + Cont. 88.65
Table 5: Performance comparison of the crash prediction models using complexity-infused features from different encoders. The results are reported in terms of accuracy (%). Upward arrows next to the values indicate that higher accuracy is preferable.
The models using different encoders consistently achieved higher accuracy as the number of input features increased. The encoder with 32 hidden neurons and a continuous output consistently delivered the highest accuracy, regardless of the input feature set. These results further validate the proposed encoder structure’s effectiveness in extracting hidden context from the existing semantic, driving, and contextual features through the complexity index.
5 Conclusions and Limitations
In this paper, we presented a two-stage framework for extracting hidden context from semantic, driving, and contextual features, and for predicting crash likelihood by incorporating these hidden context features into the original feature sets. This approach addresses the challenges of predicting crashes in real-world and naturalistic settings. To our knowledge, this is the first model to integrate all scene-related and driving-related features and link them to a complexity index obtained from the LLMs model. Our experiments revealed that the framework can accurately predict crash likelihood, achieving an accuracy of $90.15\%$. The data generated by LLMs provided better predictive capability than that obtained from MTurk, and different encoder structures consistently outperformed models without complexity-infused features. This solution has the potential to inform the development of warning systems and driving monitoring systems in human-machine interfaces, improve safety in autonomous driving, and enhance human-AI collaboration.
In the future, we aim to expand the dataset by incorporating more diverse driving environments and conditions, such as adverse weather and complex urban settings, to further improve the generalizability of the model. In addition, we intend to experiment with more advanced encoder architectures to further refine the extraction of hidden context features. Finally, we plan to investigate the integration of real-time data, enabling dynamic crash likelihood predictions for real-world applications.